# Deep learning framework for action prediction reveals multi-timescale locomotor control

### v4

@@ Results
The timing at which the swing foot begins to contain significant information regarding foot placement (“breakpoint” timing, Figure 5C) provides further insight into the swing foot prediction dynamics. During treadmill walking, lateral foot placement prediction (FP) timing occurs significantly later than the swing foot initiation (Wilcoxon signed-rank test, one-sided, $p<0.05$; Figure 4B), indicating that the swing foot does not contribute substantial information to foot placement immediately after the swing begins. However, no significant difference is observed between swing timing and fore-aft FP timing. This finding is consistent with overground walking on flat terrain, where lateral FP timing occurs later than swing timing, while fore-aft FP timing does not exhibit such a delay (Figure 4C). As terrain complexity increases during overground walking, both lateral and fore-aft FP timing are delayed relative to swing timing, perhaps reflecting the additional planning demands imposed by rougher terrains. A similar delay is observed during treadmill running, where both lateral and fore-aft FP timing occur much later than swing initiation. This is likely due to the earlier swing initiation in running driven by the flight phase, while the swing foot does not contribute significant information about foot placement until later in the gait cycle (Figure 4C). These findings highlight how swing foot prediction dynamics are adapted to environmental and task-specific demands, and that swing initiation does not necessarily coincide with the point at which the swing foot begins to contain significant information about foot placement.
Figure 4: Swing foot dynamics predict future foot placement mid-swing. A Timing of when the swing foot starts to swing across different contexts. B, C Comparison of the timing when the swing foot starts to swing with the timing of when the swing foot begins to provide significant information about foot placement, detected using the average slope method (Figure 5C), during treadmill (B) and overground (C) movements. ∗∗ for $p<0.01$; $n.s.$ for not significant.
@@ Discussion
In this study, we present a data-driven framework to characterize real-world motor behaviors by how much and when different input modalities predict future actions. We incorporate trial embedding information into the framework in an architecture-dependent manner, which enables the models to separate trial-specific explained variance from the variance explained by the input modality. By employing deep learning models that incorporate the temporal history of state estimates, we outperform simpler models in non-stationary tasks for predicting future foot placements. We compare the predictive power of various hypothesized input modalities (e.g., center-of-mass kinematics, full-body kinematics, and gaze fixations) against an autoregressive baseline. This baseline comparison reduced the impact of inter-trial variability, allowing us to isolate and interpret the context- and modality-dependent effects more clearly. We quantify the contribution of each modality to foot placement through their distinct prediction timescales. We discover a tradeoff between long- and short-timescale predictions across contexts; in more uncertain contexts like rough terrain, we find a reduced reliance on long timescale predictions. Finally, rather than assuming that foot placement is initiated at the very start of the swing cycle, our method helps identify the swing phase at which the foot begins to carry significant predictive information about the foot placement location. Together, these findings highlight multi-timescale control strategies in real world locomotion, revealing how they are shaped by the task context and the input modality. Our framework can be extended to analyze other input modalities and motor actions to characterize real world motor behavior.
@@
Comparing different model architectures in their ability to predict foot placement.
@@
The model performances across tasks and input modalities (Figure 1) highlight the strengths and limitations of various network architectures. During laboratory-constrained treadmill locomotion, linear models demonstrate performance comparable to the best-performing nonlinear models. This suggests that control strategies during treadmill locomotion is relatively simple and linear [42, 43], and therefore can be effectively captured by existing models that assume fixed-timescale linear mappings [7, 2]. Interestingly, the Temporal Convolutional Network (TCN) performed worse than linear models, which can be attributed to its rigid structural design, especially its reliance on fixed-size convolutional kernels and local receptive fields. On treadmill locomotion tasks, where there is less input history-dependence, fully connected neural network (FCNN) performs well as a result of the simplicity and linearity of the mappings [44]. During overground walking, we observe that GRUs [45] and Transformers [46] perform the best, since their gating or attention mechanisms enables the models to dynamically focus on the most relevant timesteps. This implies that the task demands are more non-stationary, and there may be varying temporal dependence on input history. While existing models, which are predominantly simple and linear, provide valuable insights into foot placement control in constrained environments like treadmill walking [7, 2, 11], they fail to capture the complexity of more naturalistic locomotion such as overground walking. During overground walking, TCNs show comparable performance to other nonlinear models, as their ability to process long-term dependencies becomes more relevant [47, 48]. In contrast, the FCNN now performs the worst among nonlinear models, highlighting the benefits of architectures that specialize in time series prediction in non-stationary scenarios. This is consistent with prior work showing that models explicitly designed for sequential data often outperform generic architectures in handling temporal dependencies [49, 50, 51].
Tradeoffs between long- and short-timescale action predictions.
@@
Our study identifies long- and short-timescale control strategies hidden in real-world movement data (Figure 2A). Previous research has highlighted the importance of fast-timescale within-step corrections, such as those used for obstacle avoidance or unexpected terrain changes to enable stable locomotion [52, 53, 7], showing that these corrections have a stabilizing effect on the body. Studies that have analyzed gaze fixations, on the other hand, have identified longer-timescale control strategies [30], but have not tested whether these co-exist with or replace fast-timescale strategies. This gap is important to address, because the hierarchical motor control hypothesis posits the existence of both low-level fast timescale and high-level slow timescale processes [54, 27]. Here, we discover the existence of both gaze-predictive long timescale and body state-predictive fast timescale control strategies during natural locomotion. For walking on uneven terrain with varying complexity, we discover a context-dependent tradeoff between slow- and fast-timescale prediction (Figure 2C). We interpret this tradeoff to reflect the greater prioritization of fast-timescale strategies in more complex environments, due to the rapid changes in the environmental statistics as a function of time. This tradeoff could be explained as the outcome of an optimization process in which working-memory resources used to represent the environment are allocated efficiently [21], relying on more recent information in more complex environments. Indeed, while optimal feedback control suggests that there is more reliance on feedback in the presence of uncertainty [55, 56, 57], here we provide evidence to support this normative hypothesis in a real world settings (Figure 2C). Our model advances the understanding of how predictive planning and reactive corrections are balanced in dynamic environments, highlighting the efficiency of the motor control system in integrating long- and short-timescale strategies to achieve smooth and adaptive movements.
Input modality-dependent timescales for motor control.
@@
By analyzing the relative predictive power of different input modalities our approach enables identification of when each modality becomes useful for planning future actions (Figure 3). Specifically, we find that full-body kinematics predict future foot placement earlier than CoM-relevant kinematics during treadmill running and overground walking. This earlier full-body prediction could be explained by the hypothesis that preparatory actions distributed across body segments may provide early cues to help modulate the CoM’s state [58]. Alternatively, earlier full-body predictions could be explained by the correlation with features not captured by the CoM state such as whole-body angular momentum and its role in foot placement prediction [59, 60]. We also found that the utilization of CoM-relevant kinematics is delayed during overground walking compared to treadmill walking. This delay likely stems from the increased variability in terrain and the demands of path planning in overground environments, which require more time to integrate information needed to determine optimal foot placement [61, 62]. Additionally, as terrain roughness increases, there is a greater reliance on gaze, as evidenced by the increasing relative predictive power of gaze fixations as an input modality [13, 22]. Previous work has demonstrated that humans use visual information for path planning on complex terrain, actively avoiding steep steps with large height changes in favor of flatter, more circuitous paths [25]. Our findings further suggest that visual inputs are primarily utilized for lateral foot placement (Figure 3E) rather than fore-aft placement (Figure 8) on uneven terrains, which aligns with previous findings that humans tend to adjust step width rather than step length to avoid obstacles [63, 64]. Future studies could build on our data-driven approach to provide insight into the control mechanisms underlying these modality-dependent timescales.
@@
Our work highlights how complex real world movement data can be combined with data-driven approaches to obtain insight into the control strategies at play. We put forth a deep learning framework for action prediction, finding that nonlinear models outperform traditional linear models by capturing the input history-dependence at play in real world contexts. Our framework provides a comprehensive and generalizable approach to compare the influence of different input modalities and contexts on the control strategies chosen. By identifying specific phases when there is more predictive ability, it provides new insight into the temporal structure of motor planning and control in high-dimensional tasks. Overall, we present a behavioral data-driven approach that future work can build on to understand context- and modality-dependent control strategies in everyday movements. Our approach can be extended to analyze large-scale data across different motor behaviors [65, 66] and species [67, 68]. Integrating the context- and modality-dependent timescales identified by our approach into physics-based models could yield more human-like simulation and clarify the underlying neuromechanics [69, 70, 71]. These data-driven models can also enable more human-like assistive robots that generalize better across environmental contexts or more directly incorporate gaze data to inform control.
@@ Methods
In this section, we present the data-driven framework for analyzing and inferring control strategies across contexts. First, we describe the datasets used in this study and how the inputs and outputs for the models were structured. Next, we introduce the linear and nonlinear models used to map the hypothesized input modalities to foot placement. We further provide details on the training and evaluation procedures for these models, and discuss the statistical tests employed to examine the prediction timescale.
@@
Our framework is versatile and can be applied to any dataset with time-series inputs and discrete output events. To demonstrate this flexibility, we utilized several existing datasets of human locomotion across a range of contexts (Figure 1A), including walking and running on both treadmills and overground, as well as across uneven terrain. The treadmill walking dataset was collected by Wang and Srinivasan [7], while the treadmill running dataset on even and uneven terrains was collected by Voloshina and colleagues [37]. The overground walking dataset, collected by Matthis and colleagues [30], includes full body kinematics of subjects walking on flat, medium, and rough terrains alongside gaze information recorded to track ground fixations during locomotion.
For all studies, we processed kinematic data by a zero-lag 4th order Butterworth lowpass filter (cutoff frequency of 6 Hz). Here, let $x$ denote the lateral direction, $y$ denote the fore-aft direction, and $z$ denote the vertical direction. For treadmill movement datasets, the fore-aft marker positions were adjusted to account for belt speed to allow direct comparisons with overground data. Formally, let $y(t)$ denote the forward position of any marker at time $t$, the adjusted fore-aft position is defined as $y^{\prime}(t)=y(t)+v\cdot t$, where $v$ is the belt speed. For overground data, the fore-aft axis was aligned with the main direction of locomotion, such that the pelvis position starts at $(0,0)$ on the x-y plane and moves towards the positive y-axis. We computed the velocity of each marker using 4th order centered finite differences of their positions. Next, we computed the heel strike timings as the time at which the fore-aft distance between the foot and the pelvis markers is maximal [72]. While other gait segmentation procedures based on velocity threshold produce similar estimates of contact detection [73], we chose to adopt the method based on the relative foot position as it was more robust to noise during overground locomotion. We segmented the data into individual gait cycles and temporally interpolated them into 20 equally spaced gait phases. We implemented several tests to detect and discard abnormal gait cycles. These tests verified whether the stance foot was fixated on the floor during the stance phase (walking data only), and whether the timing of each heel strike was coherent with neighboring heel strikes. Overall, less than 0.1% of the steps were discarded for treadmill walking and running, while less than 2% of the steps were discarded for the overground locomotion dataset.
@@
Foot placement control depends on the history of input modalities, such as postural or gaze information. In this study, we leveraged deep learning techniques to evaluate the correlations between various input modalities and future foot placement (see Figure 1 for an overview of the modeling framework).
We hypothesized that different input modalities are integrated throughout locomotion to inform future foot placements. These modalities include the CoM-relevant kinematics, the swing-foot kinematics, the full-body kinematics, and the gaze fixations (Figure 1B). The output predicted by the models consists of the lateral and fore-aft foot placement relative to the contact location of the opposite foot at heel strike (Figure 1C). In this study, we leveraged deep learning models to evaluate the relationships between these input modalities and the future foot placement. Let $v_{i}\in\{0,1,\dots,T-1\}$ be the trial ID of data point $i$, where T is the number of trials, and $l_{i}\in\left\{0,1\right\}$ be the left-right (L/R) flag of data point $i$, representing whether the corresponding output data point is a left (0) or a right (1) heel strike. The time series capturing the input modalities is denoted by $\mathbf{S}_{i}=\left[\mathbf{s}_{i}^{(1)},\mathbf{s}_{i}^{(2)},\dots,\mathbf{s}_{i}^{(m)}\right]$, where m is the number of features, and $\mathbf{s}_{i}^{(j)}$ is the $j$-th feature of the time series. This time series starts 6 steps prior to the predicted heel strike (Figure 1D) and its end is sampled from the previous gait cycle with 20 equally spaced gait phases. This range is chosen because 6 steps is a reasonable upper bound of how much subjects plan their foot placement ahead, while 2 steps is a reasonable lower bound [7, 2, 30]. Let $\phi$ denote the gait phase relative to one gait cycle prior to heel strike (i.e. previous heel strike of the same foot). With the dataset sampled at 20 time steps per gait cycle, the length of $\mathbf{s}_{i}^{(j)}$ ranges from 41 (when time series ends at $\phi=0$) to 61 (when time series ends at $\phi=1$). We define the concatenated inputs as $X=\{x_{i}\}_{i=1}^{n}$, where $x_{i}=\{\mathbf{S}_{i},v_{i},l_{i}\}$, and seek to obtain the mapping between those inputs $X$ and the outputs $Y=\{y_{i}\}_{i=1}^{n}$, where $y_{i}=\left[f_{i}^{ML},f_{i}^{AP}\right]$ where $f_{i}^{ML}$ and $f_{i}^{AP}$ represent the mediolateral (ML) and anteroposterior (AP) foot placement, respectively.
@@
Training deep learning models typically requires large amounts of data [74], which is challenging with locomotion as most datasets only capture a few minutes of data for each trial. Therefore, it is often impractical to train deep learning models tailored to specific subjects or trials. To address this limitation, we pooled the data from all trials and train a unified model that includes trial ID as an input. By doing so, the model can learn trial-specific variations in a data-driven way while efficiently leveraging the entire dataset (Figure 1E). We first reshaped the left-right flag $l_{i}$ to match the time series input’s dimension and concatenated them, which resulted in a matrix of shape $(t,m+1)$, where $m$ is the number of features in the time series. The trial ID was embedded [75] into a vector $E$ of length $\left|E\right|=\lceil\sqrt{T}\rceil$ and integrated with the time series data in a manner that is dependent on how each neural network processes the timesteps. The embedding enables us to isolate the inter-trial variability from the context- and modality-specific variability, which are the focuses of this study. The models we used include:
@@
Nested cross-validation was employed to evaluate the models’ performance and hyperparameter tuning, ensuring an unbiased estimate of model performance and preventing data leakage [78]. The dataset was initially split into five equal folds. In each iteration of the outer loop, 80% of the data was allocated for training and validation, while the remaining 20% was held out as the testing set. The training/validation set was further subdivided into five inner folds. Four inner folds were used to train the model, and the remaining fold was used for early stopping, with a patience of 50 epochs. This process was repeated for every combination of hyperparameters, with parameter updates performed using the ADAM optimizer [79] and a learning rate of 0.001. We used mean squared error (MSE) as the loss function. After each training session, the model’s performance was evaluated on the outer testing fold. The outer loop was repeated five times, ensuring that each data point in the original dataset was used once for testing and four times for training/validation.
@@
Optimal hyperparameters for each modality and dataset were first identified using grid search (or random search for large hyperparameter spaces exceeding 100 configurations; see Appendix for details). These hyperparameters were then used in the nested cross-validation process. Aggregated predictions from all outer folds, covering 100% of the data points, were used to compute the explained variance ($R^{2}$) of foot placement predicted by the input modality. Under this nested cross-validation framework, it is possible, though unlikely, to obtain negative $R^{2}$ values, which occurs when the model fails to generalize to the testing dataset, performing worse than simply predicting the average foot placements (which would result in $R^{2}=0$). In such cases, the model’s predictive power is minimal or nonexistent. To prevent overfitting, we used early stopping and dropout when training the neural networks, whereas $L2$ regularization is used to prevent overfitting to historical data or high-dimensional input space for linear models. Since neural networks are more complex than linear models, they inherently exhibit lower bias and higher variance. To mitigate this higher variance of neural networks and to better interpret model behavior across gait phases, we smooth the $R^{2}$ curves using local regression (LOWESS) and cubic spline.
@@
To evaluate model performance, we assigned a score to each model based on the Root Mean Squared Error (RMSE). As the time window increases, the input contains accumulative information about the foot placement thereby enhancing the models’ predictive power. We used the RMSE curve achieved by the swing-foot kinematics as a baseline, reflecting the predictive information inherently contained in the swing foot dynamics [80]. We define the model score as the RMSE as defined in next section). The normalized model score is then computed relative to the model with the highest score. Formally, let $\text{RMSE}(m,\phi)$ denote the RMSE of model $m$ at gait phase $\phi$ and let $c$ denote the critical gait phase. We define the model score $s_{m}$:

### v5

@@ Results
Swing foot control is an essential strategy for achieving stable locomotion in varying terrains and tasks. In prior work, it is implicitly assumed that swing control during walking occurs simultaneously with the timing of swing initiation [11, 12, 30]. However, our data-driven framework reveals that swing foot initiation timing is not necessarily the same as when the swing foot begins to predict its future placement. We find that the timing when the swing foot begins to contain significant information to predict its future placement, namely the “breakpoint” timing, is different from the swing initiation timing. For the purpose of this comparison, the timing of swing initiation is defined as when the swing foot’s velocity exceeds 5% of its peak velocity. We find that, during treadmill and overground walking on flat terrain, the lateral foot placement prediction (FP) timing occurs significantly later than swing foot initiation (Wilcoxon signed-rank test, one-sided, $p<0.05$; Figure 4). However, we do not find this to be true for fore-aft FP timing. On the other hand, for overground walking on uneven terrain, both the lateral and fore-aft FP timing were delayed relative to swing initiation (Figure 4). Similarly, during treadmill running, where swing initiation occurs earlier due to the existence of a flight phase, both lateral and fore-aft FP timing occur after swing initiation (Figure 4). This further highlights that while the swing foot begins its motion early, it need not contribute information about foot placement until later in the gait cycle [31, 32]. Taken together, our findings reveal how swing foot kinematics are adapted to environmental and task-specific demands, and that swing initiation does not necessarily coincide with the point at which the swing foot begins to predict its future placement. This delay could reflect the additional planning demands of navigating complex environments.
Figure 4: Swing foot kinematics predict future foot placement mid-swing. Comparison of the timing when the swing phase begins to when the swing foot begins to predict its future placement, detected using the methods outlined in Figure 6. Comparisons are shown for all tasks and environmental contexts including treadmill walking, treadmill running and overground walking across terrain variations. ∗∗ indicates $p<0.01$; $n.s.$ indicates not significant.
Relative predictive power reveals different control timescales across input modalities
Human locomotion requires the dynamic integration of multiple sensory inputs to guide control. Here, we quantify the timescales at which different input modalities are used for foot placement control, including center of mass (CoM)-relevant kinematics, full-body kinematics, and gaze fixations. To do this, we trained our data-driven framework to predict future foot placement based on time series data from each modality. As the input time window increases, the predictive power of our models also increases (Figure 5). We quantified the control timescale for each input by analyzing its relative predictive power ($\Delta R^{2}$) i.e. the improvement over a baseline model trained only on the foot state itself (Figure 5, inset).
We found that full-body kinematics began to outperform the baseline model earlier than CoM-relevant kinematics. Despite including significantly more body markers (29 vs. 1), full-body kinematics only explained an additional 6–14% of foot placement variance. This highlights the primary role of CoM-relevant kinematics in locomotor control (Figure 5). During walking on more complex terrain, gaze was a more accurate predictor of foot placements than body states. Gaze-based predictions occurred earlier than both CoM-relevant and full-body kinematics for lateral (Figure 5A) but not for fore-aft foot placements (Figure 9C). This indicates that in challenging environments, humans use visual information to inform lateral foot placement earlier than information from the body states. Across environmental contexts and input modalities, lateral foot placement prediction timescales were earlier than the fore-aft timescales (Figure 9A). This suggests a prioritization of lateral stability during locomotion, which is consistent with previous findings [11, 33, 34]. In the case of treadmill running, the timescale of lateral foot placement prediction was not significantly different between even and uneven terrain. However, predictions for fore-aft foot placement occurred earlier on uneven terrain (Figure 9B). This discrepancy is likely because the uneven terrain running data only included fore-aft perturbations [35].
Figure 5: Relative predictive power reveals modality-dependent control timescales. A Relative predictive power ($\Delta\,R^{2}$, median and interquartile range) of CoM-relevant kinematics, full-body kinematics, and gaze fixations for overground walking on different terrain roughness. The highlighted interval begins when $\Delta\,R^{2}$ exceeds 5% (Wilcoxon, one-sided), and ends at the peak $\Delta\,R^{2}$. The shaded area represents the swing phase. B Relative predictive power of CoM-relevant kinematics for treadmill walking, and CoM-relevant kinematics and full-body kinematics for treadmill running on even and uneven terrains.
@@ Discussion
In this study, we discover that the timescale of locomotor control strategies depend on the environmental context as well as the sensory input modality. We present a data-driven framework to predict future motor actions from time-varying input modalities across real-world contexts. By employing deep learning architectures that incorporate the temporal history of state estimates, we outperform previous models in predicting future foot placements during non-stationary tasks. We compare the predictive power of various input modalities (e.g., center-of-mass kinematics, full-body kinematics, and gaze fixations) against a baseline model. Using this framework, we reveal a tradeoff between fast- and slow-timescale predictions across contexts, where humans rely more on fast timescale predictions in uncertain contexts such as rough terrain. We quantify the contribution of each modality to foot placement through their distinct prediction timescales, finding that gaze predicts future foot placements before the body states, and the full-body states predict foot placement before the center-of-mass states. Finally, we identify the swing phase at which the foot begins to carry significant predictive information about the future foot placement, which does not necessarily correspond to the start of the swing phase. The models we developed can be integrated with locomotor rehabilitation technologies and simulations to improve their generality across environmental contexts. This framework can be extended to other input modalities and motor actions to characterize real world motor behavior.
@@
Comparing different model architectures in their ability to capture locomotor control.
@@
The model performances across tasks and input modalities (Figure 1) highlight the strengths and limitations of different network architectures. During laboratory-constrained treadmill locomotion, linear models demonstrate performance comparable to the best-performing nonlinear models. This suggests that control strategies during treadmill locomotion are relatively simple and low dimensional [36, 37], and therefore can be effectively captured by previous models that assume fixed-timescale linear mappings [11, 12]. Interestingly, the Temporal Convolutional Network (TCN) performed worse than linear models, which can be attributed to its rigid structural design, or more specifically, its reliance on fixed-size convolutional kernels and local receptive fields, making it more prone to overfitting to noise. During treadmill locomotion tasks, where there is less input history-dependence, fully connected neural networks (FCNN) perform well as a result of the simplicity and linearity of the mappings [38]. During overground walking, we observe that GRUs [39] and Transformers [40] perform the best, since their gating or attention mechanisms enable the models to dynamically focus on the most relevant timesteps. This implies that when the task demands are more non-stationary i.e. when the statistics of the environment are time-varying, this will influence the timescale of the input history dependence for control. While previous models provide valuable insights into foot placement control in constrained environments like treadmill walking [11, 12, 30], they fail to capture the complexity of more naturalistic locomotion such as overground walking. During overground walking, TCNs show comparable performance to other nonlinear models as their ability to process long-term dependencies becomes more relevant [41, 42]. In contrast, the FCNN now performs the worst among nonlinear models, highlighting the benefits of architectures that specialize in time series prediction in non-stationary scenarios. This is consistent with machine learning research in other domains which show that models explicitly designed for sequential data often outperform generic architectures in handling temporal dependencies [43, 44, 45].
Tradeoffs between fast- and slow-timescale action predictions.
@@
Our study identifies fast- and slow-timescale control strategies hidden in real-world movement data (Figure 3A). Previous research has highlighted the importance of fast-timescale within-step corrections, such as those used for obstacle avoidance or unexpected terrain changes to enable stable locomotion [46, 47, 11], showing that these corrections have a stabilizing effect on the body. Studies that have analyzed gaze fixations, on the other hand, have identified slower-timescale control strategies [5], but have not tested whether these co-exist with or replace fast-timescale strategies. This gap is important to address, because the hierarchical motor control hypothesis posits the existence of both low-level fast timescale and high-level slow timescale processes [48, 10]. Here, we discover the existence of both gaze-predictive slow timescale and body state-predictive fast timescale control strategies during natural locomotion. For walking on uneven terrain with varying complexity, we discover a context-dependent tradeoff between slow- and fast-timescale prediction (Figure 3C). We interpret this tradeoff to reflect the greater prioritization of fast-timescale strategies in more complex environments, due to the rapid changes in the environmental statistics as a function of time. This tradeoff could be explained as the outcome of an optimization process in which working-memory resources used to represent the environment are allocated efficiently [25], relying on more recent information in more complex environments. Indeed, while optimal feedback control suggests that there is more reliance on feedback in the presence of uncertainty [49, 50, 51], here we provide evidence that could help extend this theory to multi-timescale control (Figure 3C). Our model advances our understanding of how predictive planning and reactive corrections are combined in dynamic environments, characterizing how the sensorimotor control system integrates fast- and slow-timescale strategies to achieve adaptive movements.
Input modality-dependent timescales for locomotor control.
@@
By analyzing the relative predictive power of different input modalities our approach identifies when each of them becomes useful for controlling future actions (Figure 5). Specifically, we find that full-body kinematics predict future foot placement earlier than CoM-relevant kinematics during treadmill running and overground walking. This earlier full-body prediction could be explained by the hypothesis that preparatory actions distributed across body segments may provide early cues to help modulate the CoM’s state [52]. Alternatively, earlier full-body predictions could be explained by the correlation with features not captured by the CoM state such as whole-body angular momentum and its role in foot placement prediction [53, 54]. We also found that the utilization of CoM-relevant kinematics is delayed during overground walking compared to treadmill walking. This delay likely stems from the increased variability in terrain and the demands of path planning in overground environments, which require more time to integrate information needed to determine optimal foot placement [55, 56]. Additionally, as terrain roughness increases, there is a greater reliance on gaze, as evidenced by the increasing relative predictive power of gaze fixations as an input modality [27, 26]. Previous work has demonstrated that humans use visual information for path planning on complex terrain, actively avoiding steep steps with large height changes in favor of flatter, more circuitous paths [6]. Our findings further suggest that gaze information primarily predicts lateral (Figure 5A) rather than fore-aft foot placement (Figure 9) on uneven terrains, which aligns with previous findings that humans tend to adjust step width rather than step length to avoid obstacles [57, 58].
@@
Our work provides a data-driven framework for understanding the control timescales at play during locomotion by combining real-world movement data with machine learning approaches. By predicting future actions from past input states, our framework characterizes these control timescales and demonstrates that nonlinear models consistently outperform traditional linear models across various environmental contexts. This approach is comprehensive and generalizable, allowing for a systematic comparison of how different input modalities and contexts influence control. By pinpointing the timescales at which past states best predict future actions, our work offers data-driven insights into the temporal structure of motor planning and control in complex, high-dimensional tasks. Our framework can be extended in future work to analyze large-scale datasets from different motor behaviors [59, 60] and species [61, 62]. The context- and modality-dependent timescales identified by our approach can also be integrated into physics-based simulations of human movement. This integration could lead to more human-like predictive simulations across different environmental contexts and help test hypotheses in-silico about the sensory and neural basis of the prediction timescales we discover (Figure 1A) [63, 64, 65]. Additionally, these data-driven models can be integrated into rehabilitation technologies to improve their applicability to everyday settings. They can be integrated into real-time biofeedback systems [66, 67] to provide predictive feedback about body states, helping to determine which states to relay and at what timescale. Similarly, our models can be combined with wearable robots , enabling them to simultaneously use both gaze [68] and body state [69] signals to inform planned assistive profiles.
@@ Methods
In this section, we present our data-driven framework for analyzing and inferring control timescales across environmental contexts and input modalities. First, we describe the datasets used in this study and how the inputs and outputs for the models were structured. Next, we introduce the model architectures used to predict foot placement from the different input modalities. We further provide details on the training and evaluation procedures for these models, and discuss the statistical tests employed to examine the timescales.
@@
Our framework is versatile and can be applied to any dataset with time-series inputs and discrete output events. To demonstrate this flexibility, we utilized several existing datasets of human locomotion across a range of contexts (Figure 2A), including walking and running on both treadmills and overground, as well as across uneven terrain. The treadmill walking dataset was collected by Wang and Srinivasan [11], while the treadmill running dataset on even and uneven terrains was collected by Voloshina and colleagues [35]. The overground walking dataset, collected by Matthis and colleagues [5], includes full body kinematics of individuals walking on flat, medium, and rough terrains alongside gaze information recorded to track ground fixations during locomotion.
For all studies, we processed kinematic data using a zero-lag 4th order Butterworth low-pass filter (cutoff frequency of 6 Hz). Here, let $x$ denote the lateral direction, $y$ denote the fore-aft direction, and $z$ denote the vertical direction. For treadmill movement datasets, we adjusted the fore-aft marker positions to account for belt speed to allow direct comparisons with overground data. Formally, let $y(t)$ denote the forward position of any marker at time $t$, the adjusted fore-aft position is defined as $y^{\prime}(t)=y(t)+v\cdot t$, where $v$ is the belt speed. For overground data, the fore-aft axis was aligned with the main direction of locomotion, such that the pelvis position starts at $(0,0)$ on the x-y plane and moved towards the positive y-axis. We computed the velocity of each marker using 4th-order centered finite differences of their positions. Next, we computed the heel-strike timings as the time at which the fore-aft distance between the foot and the pelvis markers is maximal [70]. While other gait segmentation procedures based on velocity threshold produce similar estimates of contact detection [71], we chose to adopt the method based on the relative foot position as it was more robust to noise during overground locomotion. We segmented the data into individual gait cycles and temporally interpolated them into 20 equally spaced gait phases. We implemented tests to detect and discard abnormal gait cycles. These tests verified whether the stance foot was on the ground during the stance phase (walking data only), and whether the timing of each heel-strike was reasonable relative to neighboring heel-strikes. Overall, we discarded fewer than 0.1% of the steps for treadmill walking and running, and fewer than 2% of the steps for the overground locomotion.
@@
Foot placement control can depend on a history of input modalities, such as postural or gaze information. In this study, we developed a data-driven framework to evaluate the ability of various input modalities to predict future foot placement (see Figure 2 for an overview of the modeling framework).
We would like to analyze whether a history of inputs are integrated throughout locomotion to inform future foot placements. These inputs can include CoM-relevant kinematics, swing foot kinematics, full-body kinematics, and gaze fixations (Figure 2B). The output predicted by the models consists of lateral and fore-aft foot placement relative to the contact location of the opposite foot at heel-strike (Figure 2C). In this study, we leveraged data-driven models to evaluate the relationships between these input modalities and future foot placement. Let $v_{i}\in\{0,1,\dots,T-1\}$ be the trial ID of data point $i$, where T is the number of trials, and $l_{i}\in\left\{0,1\right\}$ be the left-right (L/R) flag of data point $i$, representing whether the corresponding output data point is a left (0) or a right (1) heel-strike. The time series capturing the input modalities is denoted by $\mathbf{S}_{i}=\left[\mathbf{s}_{i}^{(1)},\mathbf{s}_{i}^{(2)},\dots,\mathbf{s}_{i}^{(m)}\right]$, where m is the number of features, and $\mathbf{s}_{i}^{(j)}$ is the $j$-th feature of the time series. This time series starts 6 steps prior to the predicted heel-strike (Figure 2D) and its end is sampled from the previous gait cycle with 20 equally spaced gait phases. This range is chosen because 6 steps is a reasonable upper bound for the amount of information that can be used to plan future foot placement, while 2 steps is a reasonable lower bound [11, 12, 5]. Let $\phi$ denote the gait phase relative to one gait cycle prior to heel-strike (i.e. previous heel-strike of the same foot). With the dataset sampled at 20 timesteps per gait cycle, the length of $\mathbf{s}_{i}^{(j)}$ ranges from 41 (when time series ends at $\phi=0$) to 61 (when time series ends at $\phi=1$). We define the concatenated inputs as $X=\{x_{i}\}_{i=1}^{n}$, where $x_{i}=\{\mathbf{S}_{i},v_{i},l_{i}\}$, and seek to obtain the mapping between those inputs $X$ and the outputs $Y=\{y_{i}\}_{i=1}^{n}$, where $y_{i}=\left[f_{i}^{ML},f_{i}^{AP}\right]$ where $f_{i}^{ML}$ and $f_{i}^{AP}$ represent the mediolateral (ML) and anteroposterior (AP) foot placement, respectively.
@@
Training deep neural networks typically requires large amounts of data [72], which is challenging for human locomotion in everyday environments as most datasets only capture a few minutes of data for each trial. Therefore, it is often impractical to train deep learning models tailored to specific individuals or trials. To address this limitation, we pooled the data from all trials for a given environmental context and trained a single model that includes trial ID as an input. By doing so, the model can learn trial-specific variations in a data-driven way while efficiently leveraging the entire dataset (Figure 2E). We first reshaped the left-right flag $l_{i}$ to match the time series input’s dimension and concatenated them, which resulted in a matrix of shape $(t,m+1)$, where $m$ is the number of features in the time series. The trial ID was embedded [73] into a vector $E$ of length $\left|E\right|=\lceil\sqrt{T}\rceil$ and integrated with the time series data in a manner that is dependent on how each neural network processes the timesteps. The embedding enables us to isolate the inter-trial variability from the context- and modality-specific variability, which are the focuses of this study. The models we used include:
@@
We employed nested cross-validation to evaluate the models’ performance and hyperparameter tuning, ensuring an unbiased estimate of model performance and preventing data leakage [76]. We initially split the dataset into five equal folds; in each iteration of the outer loop, 80% of the data was allocated for training and validation, while the remaining 20% was held out as the testing set. The training/validation set was further subdivided into five inner folds. Four inner folds were used to train the model, and the remaining fold was used for early stopping, with a patience of 50 epochs. This process was repeated for every combination of hyperparameters, with parameter updates performed using the ADAM optimizer [77] and a learning rate of 0.001. We used mean squared error (MSE) as the loss function. After each training session, the model’s performance was evaluated on the outer testing fold. The outer loop was repeated five times, ensuring that each data point in the original dataset was used once for testing and four times for training/validation.
@@
We first identified optimal hyperparameters using grid search (or random search for large hyperparameter spaces exceeding 100 configurations; see Appendix for details). These hyperparameters were then used in the nested cross-validation process. Aggregated predictions from all outer folds, covering 100% of the data points, were used to compute the explained variance ($R^{2}$) of foot placement predicted by the input modality. Under this nested cross-validation framework, it is possible, though unlikely, to obtain negative $R^{2}$ values, which occurs when the model fails to generalize to the testing dataset, performing worse than simply predicting the average foot placements (which would result in $R^{2}=0$). In such cases, the model’s predictive power is minimal or nonexistent. To prevent overfitting, we used early stopping and dropout when training the neural networks, whereas $L2$ regularization is used to prevent overfitting to historical data or high-dimensional input space for linear models. Since neural networks are more complex than linear models, they inherently exhibit lower bias and higher variance. To mitigate this higher variance of neural networks and to better interpret model behavior across gait phases, we smooth the $R^{2}$ curves using local regression (LOWESS[78]) and cubic spline.
@@
To evaluate model performance, we assigned a score to each model based on the Root Mean Squared Error (RMSE). As the time window increases, the input contains accumulative information about the foot placement thereby enhancing the models’ predictive power. We used the RMSE curve achieved by the swing foot kinematics as a baseline, reflecting the predictive information for future foot placement inherently contained in the swing foot dynamics [79]. We define the model score as the RMSE as defined in the next section. The normalized model score is then computed relative to the model with the highest score. Formally, let $\text{RMSE}(m,\phi)$ denote the RMSE of model $m$ at gait phase $\phi$ and let $c$ denote the gait phase at the end of the input window. We define the model score $s_{m}$: